### v1

# Understanding and predicting adsorption energetics on monolayer transition metal dichalcogenides

Abstract
Two-dimensional (2D) transition metal dichalcogenides (TMDs) have recently been shown to demonstrate non-volatile resistive switching (NVRS), offering significant advantages such as high-density integration and low energy consumption due to their atomic-scale thinness. In this study, we focus on the adsorption and desorption of metal adatoms, which can modulate the electrical resistivity by several orders of magnitude. We develop material-based relationships of the adsorption energy with electronic and atomic structure descriptors by examining the effects of various transition-metal adsorbates on the surface of TMDs. Our results reveal that adsorption energies of transition metals exhibit consistent trends across different TMDs (MoS2, MoSe2, WS2, WSe2) and can be explained using simple descriptors of the atomic and electronic structure. We propose several models to describe this adsorption process, providing a deeper understanding of a crucial step in the resistive switching mechanism based on formation and dissolution of point defects. Finally, we connect our computed adsorption energies to the switching energy. These findings will help guide rational materials selection for the development of NVRS devices using 2D TMDs.
I Introduction
Diversification of memory and storage technologies and improvements of 10$\times$-100$\times$ in capacity and energy efficiency are required, necessitating novel materials platforms [1, 2]. The drive towards smaller and more energy-efficient electronic devices has heightened interest in two-dimensional (2D) materials  [3, 4, 5, 6]. This interest stems from the unique electronic, optical, and scalability properties of the low-dimensional materials [7, 8, 9, 10, 11]. Two-dimensional materials exhibiting NVRS [12, 13, 14] have demonstrated high yields at wafer-scale [15], low cycle-to-cycle and device-to-device variability [16], and fast switching times [17, 18], making them competitive with existing and novel memory and storage technologies and neuromorphic computing applications [19]. As a result, resistive switching based on atomically-thin 2D materials, including the transition metal dichalcogenides (TMDs) [20, 21, 22, 23], are increasingly viewed as promising materials platforms to address the pressing challenges of achieving higher density integration and enhanced energy efficiency in next-generation computing technologies.
How the switching characteristics of memristors depend on the switching and electrode materials is not fully understood. In lateral configurations, where electrodes are placed on the same plane as the TMD layer, Sangwan et al. [24] demonstrated that gate-tunable memristive behavior in monolayer MoS2222 devices allows field-control of switching through electrostatic gating and operates via migration of grain boundaries via bias sweeps in the 10s of volts. For vertical memristors, with electrodes positioned above and below the TMD layer, exhibit distinct out-of-plane switching dynamics. Xu et al. [25] reported vertically stacked memristors using graphene/MoS2/graphene structures, which provided excellent flexibility and endurance and switching voltages in the 100 milliVolt range. Kumar et al. [26] further explored vertically aligned MoS2 layers with metal electrodes, demonstrating that this configuration leads to enhanced on/off ratios and retention times due to improved layer-to-layer coupling. Previous studies have achieved artificial synpases with combining different TMD layers or other 2D materials to form junctions,  [27, 28] but these studies focused on imitating switching characteristics to biological synapses and device integration rather than a materials understanding.
Several mechanisms for resistive switching have also been proposed, including due to the migration of extended defects[24], the generation of vacancies [13], the formation of a conductive filament or bridge [25, 29, 30, 31], the migration of defect ions originating from the electrode [32], and relatedly, the adsorption and desorption of a metal adatom defect on a pre-existing anion vacancy [33, 14, 34, 35, 36]. Thus far, there is a huge variability in reported switching characteristics in low-dimensional materials [5, 37, 38]. For example, MoS2 with gold electrodes experimentally has been reported to have a higher switching ON/OFF ratio compared to WSe2 [14] but the opposite trend is predicted from theory [36], highlighting a critical need to understand the switching and electrode materials and their defect chemistry. Indeed, the plethora of NVRS mechanisms simultaneously highlights the prevalence of resistive switching in low-dimensional materials and the gaps in knowledge of the atomic underpinnings in resistive switching mechanisms at the material level.
A materials-based understanding will help pave the way for exploring other transition metals, provide broader insights into optimizing memristive performance, and help in identifying new electrode and switching material combinations. Several important computational studies have been made to understand defects in 2D materials and in resistive switching, including the generation [39, 40, 41, 42] and analyses [43] of several databases. However, these databases and studies do not yet cover the space of defects relevant to resistive switching, e.g., the complex between an impurity and existing native point defect. Furthermore, to the best of our knowledge, very few limited systematic relationships have been determined for understanding the energetics of defects in low-dimensional materials. In this study, we demonstrate that insights from electrochemistry can be directly applied to understand the defect energetics of materials relevant to resistive switching applications.
Figure 1: (a) Periodic table highlighting the elements studied in this work. Adatoms investigated for adsorption on transition metal dichalcogenides (TMDs) are marked in green. The transition metal (M) and chalcogen (X) elements in the MX2 configuration are highlighted in purple and yellow, respectively. (b) Representative current-voltage (I-V) curve demonstrating resistive switching behavior in a MoS2-based memristor device. The high resistance state (HRS) is indicated in red, corresponding to the system before adatom adsorption, while the low resistance state (LRS) is shown in black, corresponding to the state after adatom adsorption.
In this study, we utilize first-principles calculations based on density functional theory[44, 45] to better understand the defect energetics in the resistive switching mechanism of monolayer transition metal dichalcogenides (TMDs). We focus on resistive switching as mediated by the formation and dissolution of point defects, namely the adsorption and desorption of a (neutral) metal adatom that originates from the electrode material on the surface of the TMD (Figure 1). Such defects are readily expected to form and contribute to resistive switching, as observed computationally [34] and experimentally [46]. In Section III.1, we discuss trends observed in computed adsorption energies for representative monolayer TMDs (MoS2, MoS2, WS2, WSe2) in order to extract meaningful materials descriptors. In order to rationalize trends in adsorption energies, we introduce a model for predicting the adsorption energies of transition metal adsorbates on TMDs, including ones based on the d-band center model inspired by the study of catalysts in Section III.2. We then present a framework for connecting the energetics of adsorption with the switching energy, inspired by the dissociation-diffusion-adsorption model [14] in Section III.4. Finally, we discuss ramifications of the study on considerations for materials selection in novel memory and computing technologies based on resistive switching mediated by point defects (Section IV).
Ultimately, this study seeks to address several key questions: Which adsorbates can induce NVRS? What dictates the adsorption energy and the switching energy? Are there discernible trends among potential materials candidates undergoing switching or for the electrode? By systematically analyzing the adsorption energies and related trends, we aim to accelerate identification of promising materials for future memristive devices based on TMDs or other 2D materials.
II Methods
II.1 Computational Parameters
The calculations in this work were performed using the Vienna Ab initio Simulation Package (VASP.6.3.0) [47, 48, 49] with the Projector Augmented-Wave (PAW) method and the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional [50, 51]. A kinetic energy cut-off of 400 eV was used for the plane-wave basis set expansion. The Brillouin zone was sampled using a 2 $\times$ 2 mesh, and spin polarization was taken into account, as it is expected to be significant when dealing with transition metal adsorbates [52]. Convergence criteria for an energy difference of $10^{-5}$ eV for electronic steps and force difference of 10 meV/Å. For density of states calculations, a $k$-point grid of 4 $\times$ 4 (equivalent to a resolution of 0.07 Å-1) is used, resulting in less than 0.01 eV difference in Fermi energies).
To investigate the overall relationship between the structure and composition of a TMD and adatom complex, we use the adsorption energy defined by Eq. 1. Here, $E_{ads}$ is the adsorption energy, which represents the energy required for the adsorption of a single adatom onto the substrate. The terms $E^{tot}_{ads/sub}$, $E^{tot}_{sub}$, and $E^{tot}_{ads}$ denote the total energies of the structure with the adsorbate, the substrate without the adsorbate, and the chemical potential of the adatom, respectively.
$E_{ads}=E^{tot}_{ads/sub}-E^{tot}_{sub}-E^{tot}_{ads}$ (1)
Figure 2: Relaxed atomic structures of MoS2 with (a) a single sulfur vacancy, (b) gold adsorbate, (c) platinum adsorbate, (d) zirconium adsorbate. The distances between the nearest neighbor molybdenum atoms in each structure are indicated.
In general, the adsorption energy will depend on the chemical potential of the adatom, and hence depend on the bath from where the adatom originates, which is depends on experimental conditions. We explore the effect of the chemical potential ranging from the isolated adatom to that referenced to the bulk metal on the adsorption energy in Section III.1. In Eq. 1, we assume the chemical potential to be effectively that of an isolated adatom in order to focus on the adsorption energetics due to chemical and electronic differences in the adsorbate or monolayer.
To minimize spurious defect-defect interactions, a supercell size of 8 × 8 (192 atoms) was used with a vacuum size of 20 Å. The relaxed structures shown in Figure 2 illustrate the distance between adjacent molybdenum atoms near the defect site. We focus on adsorption of a metal adatom on a pre-existing anion vacancy as the representative point defect complex, as this has been shown in our prior work to be more energetically favorable compared to the adsorption of the adatom on a pristine monolayer [34]. For example, adsorption energies of Ag and Au on pristine MoS2 were calculated to be -0.59 eV and -0.64 eV, respectively, at the site directly above the molybdenum, with similar energies within 0.2 eV at other sites (-0.74 eV above the chalcogen site, -0.73 at the hollow site). In contrast, the adsorption energy at the chalcogen vacancy site was calculated to be -2.56 eV for gold. This is consistent with the dissociation-diffusion-adsorption (DDA) model [14] presented in Section III.4, where the mechanism involves the dissociation of the adatom from the electrode, diffusion across the TMD, and eventual adsorption at an anion vacancy site (see also Section III.4).
II.2 The d-band model
The d-band center model is an important theoretical framework used to understand the interaction between adsorbates and transition metal surfaces in catalysis applications [53, 54, 55, 56], including those involving transition metal dichalcogenides (TMDs) [57, 58, 59]. Originally, this model was used to study the catalytic activity of bulk transition metal surfaces with adsorbates such as hydrogen [53], nitrogen [60], and carbon dioxide [61].
We apply the d-band model to rationalize trends in adsorption energy as they relate to resistive switching applications. The d-band model is based on the Newns-Anderson model [62, 63], which correlates the energy levels of the adsorbate with the electronic structure of the substrate. Subsequent refinements to this model have enhanced its accuracy and applicability, such as the inclusion of spin [52] and the inclusion of structural information via compound descriptors [64], and we utilize aspects of these refinements in our study. This approach utilizes the energy level of the d-band center and the shift induced by the adsorbate as key descriptors for adsorption energy.
The d-band center, $\varepsilon_{d}$, is defined as the first moment of the density of states (DOS) of the $d$-electrons and can be calculated as:
[…]
where $\varepsilon$ represents the energy eigenvalues, and $n(\varepsilon)$ is the density of states of the bands with projected d-character. In our calculations, we define the d-band center over a fixed energy range that includes both regions below and above the Fermi level between the different adsorbates. This approach allows for a direct comparison of d-band centers across different adsorbates. By considering a sufficiently broad energy range that spans across the occupied and occupied bands within the band gap, the d-band center reflects the overall electronic environment of the system. Here in this study, we choose an energy range of -10 eV to 1.5 eV (referencing EF = 0 eV). This strategy enables an understanding of how the electronic structure influences adsorption energy.
In our spin-polarized calculations, the d-band centers for the spin channels are combined into a single descriptor using the method developed by Bhattacharjee et al., which employs a two-centered d-band model [52, 65]. In this model, a new descriptor called the effective d-band center is introduced, as shown in Equation 3:
$\varepsilon^{\text{eff}}=\frac{\sum_{\sigma}f_{\sigma}\varepsilon_{d\sigma}}{\sum_{\sigma}f_{\sigma}}-(\varepsilon_{d\downarrow}-\varepsilon_{d\uparrow})\mu$ (3)
Here, $f_{\sigma}$ represents the fractional occupancy of spin state $\sigma$, and $\varepsilon_{d\sigma}$ is the conventional Hammer-Nørskov d-band center for spin state $\sigma$. The final term, $\mu$, is the reduced fractional moment, defined as $\mu=\frac{f_{\uparrow}-f_{\downarrow}}{f_{\uparrow}+f_{\downarrow}}$. The first term in the equation gives the spin-averaged d-band center, where the spin-up and spin-down channels are averaged, while the second term introduces a shift in the d-band center that depends on spin polarization.
Apart from the d-band center, other parameters in the d-band model include the orbital occupation, the coupling between the electronic states of the adsorbate and substrate, and the energy associated with the orthogonalization of the states between the adsorbate and substrate. These descriptors are largely based on the electronic structure and thus suggest a generality of the d-band model beyond catalysis, extending to applications such as resistive switching based on the formation of point defects.
As a complement to the d-band center model, we also utilize crystal orbital Hamiltonian population analysis (COHP) to further study the bonding nature of the adsorbate on the various TMDs [66]. These calculations are performed using the Local-Orbital Basis Suite Towards Electronic-Structure Reconstruction (LOBSTER) package [67, 68, 69, 70, 71]. Crystal Orbital Hamiltonian Population (COHP) analysis is a method used to distinguish between bonding, non-bonding, and antibonding interactions between adjacent atoms by analyzing the projected density of states (DOS) weighted by Hamiltonian matrix elements. These elements are related to the overlap integrals between atomic orbitals, allowing for a quantitative assessment of the bonding characteristics in a crystal or molecule. Specifically, the Hamiltonian element corresponds to the interaction strength between pairs of atomic orbitals on neighboring atoms, allowing COHP to quantify the extent of bonding (stabilizing) and anti-bonding (destabilizing) character in a material’s electronic structure. This method provides a detailed understanding of the nature of chemical bonding within the material by linking the DOS with the underlying electronic interactions dictated by the Hamiltonian.
Finally, we use the Sure Independence Screening and Sparsifying Operator (SISSO) [72] to develop interpretable descriptors that leverage chemical intuition to predict adsorption energy (see Table S2 in the SI for the complete list of descriptors). SISSO is a symbolic regression method that systematically combines chemically relevant primary features—such as ionization energy, electronegativity, and atomic radii—using mathematical operations to generate a vast space of potential descriptors. By creating a large pool of candidate features and applying a sparsifying operator, SISSO identifies a minimal set of the most predictive and interpretable descriptors, effectively capturing nonlinear relationships between input features and adsorption energy. This approach ensures the resulting models are physically meaningful and avoids overfitting by selecting descriptors that generalize well across different datasets. The strength of SISSO lies in its ability to reduce the problem’s dimensionality while preserving essential physics, distilling complex material behaviors into a few key descriptors. This enhances predictive accuracy and provides significant insights into the determinants of adsorption energy. Further details on the implementation are provided in Supplementary Information (SI).
III Results
We now present results on the trends of adsorption energies and their relation to the switching energies. Results in the main text are mainly presented for MoS2. Corresponding results for MoSe2, WS2, and WSe2 are provided in the Supplementary Information (SI).
III.1 Trends in the adsorption energy
[…]
These observed trends, on first examination, are correlated with the atomic electron configuration of the adatom, since one major difference among the adsorbate-substrate pairs in this study is the variation in the valence electron configurations of the adsorbates. When a chalcogen vacancy is present, the cation in the TMD is under-coordinated. The chalcogen vacancy leads to two excess electrons, which alters the electronic environment of the transition metal. When a metal adatom is introduced onto the TMD surface, it can interact with the unsaturated Mo atom, potentially altering its electronic environment further. Metals like Zn, Cd, and Hg, which have fully filled d-orbitals and s-orbitals, exhibit the weakest adsorption energies and, in some cases, even positive adsorption energies. A positive adsorption energy in this context indicates that these metals do not readily adsorb onto the vacancy. Due to their stable electron configurations, these metals do not readily donate or accept electrons, resulting in minimal interaction with the unsaturated Mo atom. Similarly, metals such as Ag, Au, and Cu, which have nearly filled d-orbitals, also show relatively weak adsorption energies. Although Cu is not a noble metal, its nearly filled d-orbital leads to weaker interactions with the TMD surface, similar to the behavior of Ag and Au. In contrast, metals with partially filled d-orbitals, such as Fe, typically exhibit intermediate adsorption energies. The overall trend of stronger adsorption energies continues when examining the early transition metals, which have mostly unfilled d-orbitals. Thus, there is a correlation with the extent of electron filling of the adatom mediating the adsorption energy.
Figure 3: Adsorption energies (left, blue axis) of various transition metal adsorbates on a single sulfur vacancy in MoS2. Adsorbates in each plot are sorted and colored by their periodic group. The right axis (red circles) displays the Bader charge gain or loss of the adsorbate relative to its neutral atomic state, with negative values indicating charge gain and positive values indicating charge loss.
We note that the adsorption energy will vary depending on the value of the chemical potential, which will depend on experimental and device conditions. Thus, we look at two limits of the chemical potential where the adatom is referenced to an isolated adatom and to the monatomic metal bulk. For the adsorption energies in Figure 3, the chemical potential is chosen as the isolated adatom, assuming dissociation of the adatom from the electrode has already occurred. When instead using the chemical potential referenced to the monatomic metal bulk, the adsorption energies can vary up to 10.93 eV, as seen in the case of tungsten (W), which had one of the largest differences between the two reference states. This difference highlights how the choice of chemical potential can affect the strength of the adsorbate-substrate interaction. A plot showing the variation of the adsorption energy with the chemical potential referenced to that of the bulk metal is provided in Figure S5 of the SI for MoS2, MoSe2, WS2, and WSe2. Referencing the chemical potential to the bulk metal results in more positive (i.e., less exothermic) adsorption energies, and reflects a scenario where the metal adatom is part of a bulk reservoir. To focus on the differences in adsorption energy arising from the interactions between the monolayer and adsorbate, we continue our discussion using the chemical potential of the isolated adatom.
III.2 Rationalizing trends in adsorption energy
To rationalize the observed trends, we employ the d-band center model, which combines the effects of electron filling and the formation of bonding and anti-bonding orbitals between the adsorbate and the substrate (also see Section II.2. Physically, the strength of interaction between the adsorbate and substrate is related to the filling of the anti-bonding states between the re-normalized valence states of the substrate and adsorbate. Because anti-bonding states generally lie energetically above the electronic band edge, the center of the d-states is an indicator of the bond strength between adsorbate and substrate. The center of the d-band is often referenced to the Fermi level, which determines the energy level of half occupation and thus delineates the filled and unfilled electronic states. Thus, a higher in energy d-band center relative to the Fermi level is indicative of a stronger adsorption interaction, as fewer anti-bonding states are occupied. Hence, the d-band model provides a basis for explaining differences and trends in adsorption energetics. Traditionally, this model focuses on the position of the d-band center of the transition metal substrate relative to the Fermi level as the typical adsorbates are organic chemical intermediates involved in catalytic reactions. By contrast, the adsorbates in this study are largely transition metal adatoms on select substrates (MoS2, MoSe2, WS2, WSe2) that show minimal difference in their structural relaxations (Figure 2). In this study, we apply the d-band center model by focusing on the d-band centers of the transition metal adsorbates instead of the TMD substrates.
Figure 4: Comparison of the d-band centers of the adsorbates, referenced to the Fermi energy, with their respective adsorption energies on MoS2. A d-band center of zero indicates alignment with the Fermi energy. The plots are categorized by specific d-block elements ((a) 3d, (b) 4d, (c) 4d, (d) all together) and color-coded by periodic groups. Linear fits with accompanying R2-values are provided as a guide the eye.
As observed from Figure 4 for MoS2, the early transition metals, such as Hf, Zr, and Ti, have d-band centers shifted closest to the Fermi level when compared to other transition metals. A similar observation can be made for MoSe2, WS2, and WSe2 (see Figures S6, S7, and S8 in the SI). According to the d-band model, this shift of the d-bands closer to the Fermi level correctly anticipates a stronger adsorption energy, as the anti-bonding states would be minimally occupied. In contrast, the late transition metals, such as Cd, Hg, Zn, Au, Cu, and Ag, have d-band centers shifted farther from the Fermi level, indicating lower energy levels for the anti-bonding orbitals, leading to weaker adsorption energy. Consequently, the d-band centers for the re-normalized electronic states of the middle transition metals are shifted by an intermediate amount, resulting in intermediate adsorption energies and strengths. By comparing the shifts of the d-band centers of different adsorbates on the surfaces of TMDs, we can directly correlate the electronic structure of the adsorbates with their adsorption energies. This approach provides a clear understanding of the trends observed across different transition metal adsorbates, highlighting the utility of the d-band center model in rationalizing trends in adsorption properties in resistive switching contexts [75].
Interestingly, we observe that by separating the transition metal adsorbates into 3d, 4d, and 5d metals, there is a more robust correlation between the computed adsorption energy and the d-band center (Figure 4). We hypothesize this to be attributed to the different (increasing) spatial extents of the 3d, 4d, and 5d transition metals, which play a role in the extent of renormalization (i.e., hybridization) between the adsorbate and surface. For example, the 3d transition metals show the lowest correlation between adsorption energy and the d-band center. In contrast, the 4d and 5d orbitals are more spatially extended and exhibit a more linear relationship between the d-band center and adsorption energy, with $R^{2}$ values of 0.78 and 0.95, respectively, for MoS2 (see also Figures S6, S7, and S8 in the SI). For the 3d orbitals, the relationship is less linear, with an $R^{2}$ value of 0.69.
We next turn to crystal orbital Hamilton population (COHP) analysis. COHP analysis provides insights into the degree of bonding and anti-bonding character in the interactions between adsorbates and substrates, helping to further understand the nature of adsorption. By decomposing the electronic density of states (DOS) into its anti-bonding and bonding contributions as a function of energy we obtain the same overall trend as seen using the d-band center model. Using the plots shown in Figure 5, the percentage contribution of anti-bonding interactions can be determined by integrating the Crystal Orbital Hamiltonian Population (COHP) curve for energies below the Fermi level.
Figure 5: Crystal Orbital Hamiltonian Population (COHP) plots of representative transition metal adsorbates from groups 11 (Au), 9 (Co), 6 (Cr), 4 (Zr) on MoS2 are presented. The blue plots show the interactions between the transition metal adsorbate and the molybdenum adsorption site in the monolayer, while the orange plots show the interactions between the transition metal adsorbate and the surrounding sulfur in the monolayer. An estimate of the percent anti-bonding character up to the Fermi level is computed for each adsorbate and corresponds to anti-bonding interactions between the adsorbate and the MoS2 atoms. The highlighted region for AuS is used to illustrate the portion of the COHP curves used to compute the percent anti-bonding character. A similar method is used for each of the different adsorbates. The gold, cobalt, chromium, and zirconium adsorbates exhibit 74.18%, 45.87%, 29.17%, and 9.62% anti-bonding character with sulfur (orange), respectively.
One may naively expect that the adsorption interactions between a TMD and a transition-metal adsorbate would be primarily governed by metal-metal interactions (for example, Mo-Au interaction). To test this whether this is true, we used our COHP analysis to decompose the overall bonding and anti-bonding interactions by species. The COHP analysis shows that Hf, Zr, and Ti have the lowest anti-bonding interactions with their nearest neighbor sulfur atoms (with percent anti-bonding contributions of 11.28%, 9.62%, and 15.39%, respectively), while Hg, Cd, and Zn exhibit the greatest anti-bonding interactions with their nearest neighbor sulfur atoms (with percent anti-bonding contributions of 84.12%, 90.0%, and 85.21%, respectively); a visualization of overall trends is presented in Figure 6. Thus, the COHP analysis shows that there is greater correlation with the anti-bonding character of chalcogen in the adsorbate interaction and the adsorption energy. We attribute this to the fact that when the adatom gains or loses charge upon adsorption, that charge is exchanged with primarily the surrounding chalcogen atoms. For example, silver adsorbed onto WSe2, the chalcogen loses a total charge of 4.34 electrons compared to a selenium vacancy in WSe2 This finding also provides physical insights as to the selection of descriptors for random forest regression of adsorption energies in the work by Frey et al [43], where it was not clear why the mean number of p-valence electrons was ranked as the highest in importance. Here, our analysis reveals that the charge transfer from the anion atoms, whose valence electrons are primarily of p-character, is a large influence on the extent of electron filling of the anti-bonding orbitals, which correlates with the strength of adsorption interactions.
Figure 6: Comparison of COHP results describing the percentage of anti-bonding interactions of the adsorbate and sulfur (ab-ADS-S) and total anti-bonding interactions in MoS2. The anti-bonding interaction percentage between the adsorbate and sulfur is correlated with adsorption energy with an R2 value of 0.70, while the combined anti-bonding character has an R2 value of 0.57.
We also conducted an analysis of the hybridization character of the adsorbates’ d-orbitals to evaluate whether differences in adsorption energy observed across different periods could be better understood. For adsorbates within the same group, despite their general similarity in electron configuration, the hybridization of the 3d, 4d, and 5d orbitals varies with the substrate. To quantify the amount of $d_{xy},d_{xz},d_{yz},d_{z^{2}}$, and $d_{x^{2}-y^{2}}$ orbital character, we integrate the projected density of states over a fixed energy range, from -2.0 eV to 0.5 eV relative to the Fermi level, to capture any mid-gap states introduced by the adsorbates; these data may be found in Figures S9 and S10 of the SI.
The variation in adsorption energies across different adsorbates have some correlation with the relative $d$-orbital contributions. For example, in the case of the group 4 adsorbates (Ti, Zr, Hf) on the MoS2 substrate, it can be observed that Ti exhibits the weakest adsorption energy of -6.13 eV. The $d$-orbital contributions of Ti are relatively evenly distributed among $d_{xy}$, $d_{yz}$, $d_{xz}$, and to a lesser extent $d_{z^{2}}$, with moderate contributions from $d_{xz}$ and $d_{yz}$. In contrast, Zr and Hf show a more pronounced contribution from the $d_{yz}$ and $d_{xz}$ orbitals, which may facilitate more effective spatial alignment with the MoS2 surface. Similarly, in group 6 metals, Mo exhibits the strongest adsorption energy of -7.78 eV on MoS2 and similarly has a larger contribution of the $d_{z^{2}}$ orbital.
Figure 7: The charge density difference between the adsorbed structure, the vS MoS2 structure, and the adsorbate plots of Nb, Y, and Zr, viewed along the b-axis and the c-axis. The yellow distribution corresponds to charge accumulation and the blue corresponds to charge depletion.
Charge Density Difference (CDD) plots provide further insight into the nature of these interactions. For instance, in the case of Nb, Y, and Zr (shown in Figure 7), which all show a higher degree of $d_{yz}$ and $d_{xz}$ contributions. The CDD plots reveal regions of charge depletion near the adsorbate and charge accumulation near the substrate. This visual evidence supports the argument that these adsorbates have ionic bonding characteristics, where the spatial alignment of the $d_{yz}$ and $d_{xz}$ orbitals of the adsorbate enables greater interaction with the substrate. Overall, the variation in adsorption energies within the same group can be partially understood to arise from differences in the hybridization and spatial orientation of specific $d$-orbitals with the substrate orbitals.
III.3 Prediction of adsorption energies
To explore the prediction of adsorption energies, we employ the Sure Independence Screening and Sparsifying Operator (SISSO) [72], a compressed sensing technique, to predict absolute adsorption energies using basic descriptors. In this context, dimensionality refers to the number of features or descriptors used to represent the data. High-dimensional datasets, where the number of potential descriptors is large, can be challenging because they increase the complexity of the model and the risk of overfitting in which the model becomes too tailored to the specific dataset and loses its ability to generalize to new data.
SISSO addresses these challenges by first reducing the dimensionality of the feature space. It does this by discarding weakly correlated features through sure independence screening (SIS), focusing the analysis on the most relevant descriptors. This reduction is crucial for managing complexity and ensuring that the model remains both robust and generalizable. Next, the sparsifying operator (SO) identifies the optimal $n$-dimensional descriptor by selecting a minimal set of features that best predict the adsorption energies, ensuring that the model is both accurate and interpretable. Unlike many machine-learning techniques, SISSO can effectively perform non-linear regression on small datasets while preserving interpretability, making it ideal for complex datasets where transparency is essential. A list of the initial set of descriptors used in the SISSO analysis is provided in the SI.
We leverage a dataset of readily available atomic properties for the substituents in each structure to construct the SISSO descriptor [72, 76]. In Section III.1, we demonstrated the significance of bonding and anti-bonding contributions to adsorption energies. This leads us to consider whether the d-band center calculation, which typically requires a detailed density of states (DOS) analysis, could be avoided. Since DOS calculations are a major computational bottleneck in predicting adsorption energy, identifying an alternative method or descriptor that bypasses this step could improve the computational efficiency of adsorption energy predictions.
To assess the performance of our SISSO-based approach, we compared the predicted adsorption energies to the values obtained via DFT calculations. Figure 8 shows the results of the best-performing compound descriptors using one- and three-dimensional descriptors. The SISSO analysis is performed for the specific case of MoS2 in Figure 8a as well as for the aggregated adsorption energies across four substrates (MoS2, MoSe2, WS2, and WSe2) in Figure 8b). When considering just the MoS2 substrate, the 1D SISSO descriptor yields an $R^{2}$ of 0.70, which improves slightly to 0.77 when increasing the dimensionality to 3D. A similar improvement is observed when comparing the 1D and 3D SISSO descriptors for all substrates considered. For the case of MoS2 alone, the 1D SISSO descriptor yields a mean absolute error (MAE) of 0.89 eV, a mean squared error (MSE) of 1.15 eV, and an R2 value of 0.71. The 3D SISSO descriptor for MoS2 improves these metrics, achieving an MAE of 0.76 eV, an MSE of 1.04 eV, and an R2 of 0.77.
When considering all substrates in combination (MoS2, MoSe2, WS2, and WSe2), the 1D SISSO descriptor results in an MAE of 0.97 eV, an MSE of 1.44 eV, and an R2 of 0.64. In comparison, the 3D SISSO descriptor for the same set achieves an improved MAE of 0.86 eV, an MSE of 1.31 eV, and an R2 of 0.70.
Figure 8: Prediction of adsorption energies from best-performing 1D and 3D SISSO descriptors for (a) MoS2 where reference DFT values are in black and SISSO values are in colored, and (b) MoS2, MoSe2, WS2, and WSe2 substrates.
Compared to other computational studies on the high-throughput calculation of defect and adsorption energies, our error metrics provide some insight in degree of error introduced when considering a minimal set of simple descriptors compared to more complex and computationally expensive non-linear regression. For instance, a study on defect formation energies in the zinc blende materials class, using random forest regression, Gaussian process regression, and neural network regression, achieved mean squared errors (MSEs) ranging from 0.94 to 1.05 eV for formation energies up to approximately 11 eV [77]. Another data-driven study on the defect formation energy of oxygen vacancies in metal oxides reported MSEs ranging between 0.21 and 0.45 eV using a suite of machine-learning regression models [78]. The approach with SISSO demonstrates promising accuracy with using only primary atomic properties—such as atomic radii and electronegativities—without the need for computationally expensive DOS calculations. This comparison highlights the efficiency of our method, such as in high-throughput studies where balancing accuracy and computational cost is crucial. It also highlights the challenges in predicting defect energetics to within chemical accuracy and emphasizes the value of a physics-informed descriptor.
Finally, we comment on the anticipated transport characteristics of the adsorbed metal adatoms, a pre-requisite for such point defects to be used in resistive switching contexts. Adsorption of gold onto sulfur vacancies in MoS2 is known to lead to up to around ten orders of magnitude of change in the normalized conductivity, including the out-of-plane component, depending on the strain present [34]. The change in conductivity is accompanied by a partially-filled defect level with predominantly Au $d$-character that imbues a degree of metallic character to the substrate. Hence, the change in conductivity can, to first order, be explained by the increase in density of states near the Fermi level [34]. When inspecting the DOS of the adsorbates studied here, we observe that 76% of the configurations tested are expected to show enhanced metallic-like behavior with the adatom adsorbed (a list of the elements may be found in Table S1 of the SI). Further investigation of the transport properties is merited and the topic of future study.
III.4 Physical model for resistive switching based on point defects
Building on the computed adsorption energies, we establish a connection between these energies and the switching energy in vertically-stacked memristors composed of monolayer TMDs. Our model of the switching energy is based on the dissociation-diffusion-adsorption (DDA) model [33, 14] (Figure S1). In the DDA model, an adatom from the electrode dissociates and then potentially diffuses across the surface before adsorbing to a pre-existing anion vacancy (Figure S1). This process involves an energy of dissociation, energy of diffusion, and energy of adsorption. This mechanism is supported through previous measurements of single-defect memristor in MoS2 based on local scanning tunneling microscopy images [46]. The energy of dissociation is determined by the chemical identity of the electrode material, and thus is assumed to be constant for a given adatom species. A full study of diffusion is outside the scope of this paper, but is the topic of on-going study. Thus, the main quantity of interest is the adsorption energy.
As discussed in Section S1 of the SI, a correlation between the measured switching energy [14] and the computed adsorption energies for the common TMDs suggests that the adsorption energy plays an important role in the energetics of resistive switching based on transition metal dichalcognides. Previous studies, such as the one by Frey et al [43], have used a similar idea of using binding energies and formation energies to identify optimal defects for applications in neuromorphic and quantum information processing. Frey et al. [43] used random forest regression to predict adsorption energies and identify potential descriptors but did not fully delve into the defect chemistry that dictated trends in adsorption energy. This study emphasizes understanding the adsorption mechanism between 2D TMDs and various transition metal adsorbates, to further elucidate how the valency, periodicity, and electronic configuration of the metal affect the adsorption energetics.
IV Discussion
We have shown that the adsorption energy is relatively insensitive to the chemical species of the substrate when considering commonly studied TMDs (MoS2, MoSe2, WS2, WSe2). Saliently, the trends in adsorption species may be explained using the d-band model, a model based on electronic structure descriptors most well-known in its application to catalysts. Hence, the switching energy is materials-dependent and that dependence can be rationalized. An earlier study on the adsorption of Li in transition metal dichachalcogenides for battery applications demonstrated that the work function can serve as a simple descriptor for explaining trends in Li adsorption[79]. In this study, we generalize this concept to adsorption energies of metal adatoms in low-dimensional materials. Thus, insights from electrochemistry and catalysis may be effectively cross-applied in resistive switching contexts, and can be applied more broadly to classes of substrates and adsorbates than previously demonstrated.
We note that these defect calculations assumed an isolated monolayer. In reality, the interactions between the TMD and the electrode at the interface are known to significantly impact the electronic properties of the TMD. For example, in our previous study on MoS2/Au heterostructures [34], the Au electrode induces a strain on the MoS2 and results in a charge transfer that alters the electronic structure and hence the resulting adsorption energies. Nevertheless, the d-band model can similarly be used to explain changes in adsorption energy due to tensile strain [34].
Additionally, in idealized models, the energetics of adsorption and desorption are often considered symmetric, meaning the energy gained via adsorption is equivalent to the energy required for desorption, and vice versa. However, in real systems where external forces like applied voltage, such as those involving resistive switching in 2D materials, this symmetry often breaks down due to surface defects, strain, and energy barriers. Thus, adsorption energies alone cannot explain the asymmetry observed in the SET/RESET I-V characteristics of measured samples [14]. We posit that this asymmetry likely requires, at minimum, the combined effects of explicitly including the presence of the electrode and an external potential to capture the full mechanism. Furthermore, while this study covers the energetics of defect formation, it does not account for the kinetics or dynamics by which defect formation occurs (e.g., diffusion processes). Each of these aspects is anticipated to play a role in the overall resistive switching characteristics of the device, thus meriting further study. While such effects are neglected here, our study provides an important baseline for understanding the intrinsic defect chemistry of the switching layer material.
Overall, we have shown with the simplifications of the dissociation-diffusion-adsorption (DDA) model[14] that performance metrics of the resistive switching, such as the switching energy and the adsorption energy, are materials dependent and can be rationalized based on materials-based descriptors of the electronic structure. For structures of 2D materials that differ from the 2H phase considered here, we anticipate incorporating structure-sensitive descriptors will further improve the model [64]. A study on the generalizability of models for adsorption energy beyond transition metal dichalcogenides or to multiple adsorption events is merited and a topic of further study.
Within and beyond resistive switching, defects in low-dimensional materials play a prominent role in next generation computing devices. Indeed, two-dimensional (2D) materials are ideal platforms for hosting quantum defects [80, 81]. An important implication of our study in rationalizing adsorption energies of point defect complexes in TMDs is the identification of electronic structure and atomic descriptors for both the substrate and electrode, which extends our ability to engineering designer defects for targeted applications. For practical device applications, we expect that the adsorption energy of transition metal adatoms plays a role in the switching energy (see also Figure S2). That is, a large (negative) adsorption energy corresponds to a greater stability of the adsorbate, which may be important for long-term operation. On the other hand, a low (near zero) adsorption energy corresponds to a low switching energy, which may contribute to more energy-efficient resistive switching.
As a practical recommendation, gold is typically chosen for its inertness to the chemical environment [75]. Our study demonstrates that this inertness extends to adsorbed adatoms, as gold exhibits a lower adsorption energy, making it favorable for resistive switching applications where energy-efficient and reversible switching is desired. Based on our calculations, we find that copper (Cu) and nickel (Ni), which are CMOS-compatible metals, have relatively low adsorption energies (-2.96 eV and -4.01 eV, respectively), which may make them suitable for switching applications with lower energy consumption. In contrast, metals like tungsten (W) and molybdenum (Mo) exhibit large adsorption energies (-7.78 eV), which may make them suitable for applications where retention of the adsorbate is needed. These findings are particularly relevant for CMOS integration, as the ability to tune adsorption energies through material selection can provide control over both the stability and switching energy of the device. Finally, we anticipate these models and descriptors to be useful in screening novel materials defects in databases [82, 41] and high-throughput calculations towards expanding our materials repertoire of low-dimensional materials for the next generation of electronic devices.
V Conclusion
We investigated the adsorption energies of various transition metal adsorbates on different monolayer TMD substrates, utilizing both computational methods and theoretical models to understand the underlying trends. We evaluated the adsorption energy of a metal adatom on a pre-existing anion vacancy for MoS2, MoSe2, WS2, and WSe2 and suggest these adsorption energies as proxy for investigating resistive switching energy. The adsorption energies displayed significant variation depending on the periodic group of the transition metal, with early transition metals (Hf, Zr, Ti) exhibiting the strongest adsorption energies, and late transition metals (Cu, Au, Ag, Zn, Cd, Hg) showing the weakest, while the middle transition metals exhibited moderate adsorption energies in between the early and late transition metals. We demonstrated these trends in the adsorption energy could be largely captured by trends in the d-band center with respect to the Fermi level, and as a descriptor can capture most of the trend in adsorption energies. COHP analysis provided further insights into the nature of bonding and anti-bonding interactions between the adsorbates and their nearest neighbor sulfur atoms in the TMDs, emphasizing the importance of the local environment surrounding the adsorption site in determining the adsorption energy. We used this understanding to perform compressed sensing on atom-based descriptors, which consequently does not require a DFT calculation. This SISSO analysis demonstrated promise in predicting absolute adsorption energies, using compound descriptors based off of the ionization energy, electronegativity, and atomic radii of the adsorbate and adsorption site.
Our findings underscore the importance of considering both the electronic structure and the local bonding environment in understanding and predicting adsorption behavior on TMD substrates. In conclusion, this study provides a framework for predicting defect energetics in TMD systems and helps to pave the way for rational materials selection in resistive switching. By integrating electronic structure analysis with techniques inspired by machine learning, we can gain deeper insights into the fundamental interactions governing adsorption and tailor materials for various technological applications in resistive switching, catalysis, sensing, and beyond.
VI Acknowledgements
This research was primarily supported by the National Science Foundation through the Center for Dynamics and Control of Materials: an NSF MRSEC under Cooperative Agreement No. DMR-2308817. The authors acknowledge the Texas Advanced Computing Center (TACC) at The University of Texas at Austin for providing computational resources that have contributed to the research results reported within this paper. D.A acknowledges the support of the National Science Foundation (NSF) award #2422934, and the Office of Naval Research (ONR) grant N00014-24-1-2080.
VII Author contributions
B.H.L: Investigation; Data collection; Formal analysis; Visualization; Writing - original draft; Writing - review & editing. J.F.: Data collection; Writing - review & editing. D.A.: Conceptualization; Writing - review & editing (supporting). W.W.: Conceptualization; Resources; Formal analysis; Investigation; Methodology; Writing - original draft; Writing - review & editing; Funding acquisition (lead); Project Administration; Supervision.
VIII Data Availability
The data supporting the findings of this study are openly available on Zenodo at DOI: 10.5281/zenodo.13892108, Ref [83]
Supplementary Information for
Understanding and predicting adsorption energetics on monolayer transition metal dichalcogenides
Brian H. Lee, Jameela Fatheema, Deji Akinwande‡, †, Wennie Wang¶, †
Texas Materials Institute, University of Texas at Austin, USA
Chandra Family Department of Electrical and Computer Engineering, University of Texas at Austin, USA
McKetta Department of Chemical Engineering, University of Texas at Austin, USA
Email: wwwennie@che.utexas.edu
S1 Connection between switching energy and adsorption energy
Figure S1: The dissociation-diffusion-adsorption (DDA) model describing the switching mechanism in a vertical MoS2 and gold memristor. The model outlines the transition between high resistance state (HRS) and low resistance state (LRS) as follows: (a) the initial state of the device, (b) dissociation of a gold adatom from the electrode, (c) weak adsorption of the gold adatom onto a pristine MoS2 site, and (d) strong adsorption of the gold adatom onto a sulfur-vacancy site.. Adapted from [14].
The switching energy in memristors can be defined as the energy required to ”switch” the material between a low-resistance state (LRS) and high-resistance state (HRS). Thus, it is expected that this switching energy is related to adsorption energies according to the dissociation-diffusion-adsorption (DDA) model (Figure S1). To demonstrate this correlation, we analyzed the I-V curves of various monolayer transition metal dichalcogenides (TMDs) exhibiting non-volatile resistive switching (NVRS), as measured by Ge et al.[14]. The switching power was estimated by numerically integrating the hysteretic I-V curves, specifically calculating the SET switching power by integrating over the SET process of the hysteresis. We report the correlation between the integrals of the I-V curves during the SET transition and adsorption energy. A stronger adsorption energy (more negative) suggests that the process is more thermodynamically favorable and is associated with a reduced power requirement to SET the device. Assuming a similar order of magnitude for the switching time across substrates (approximately tens to hundreds of nanoseconds[17, 18]), the switching power serves as an approximate measure of the switching energy.
Figure S2: Linear relationship between the computed adsorption energy of gold on various TMDs and the measured SET switching power, [14] indicating that stronger adsorption energies correlate with reduced switching power to SET the device. A linear fit is provided as a guide to eye on trend among the semiconducting materials; h-BN is omitted in the linear regression, as it is an wide band-gap insulator.
Figure S2 shows there is a correlation between the estimated measured switching power and the computed adsorption energy for various monolayer TMDs, lending credence to our model of adsorption energy as a proxy to switching power, and relatedly, switching energy. While adsorption energy is typically considered to be directly related to the RESET switching power in memristors—where the desorption and reformation of atomic or molecular states occur according to the DDA model—it is not solely related to the RESET process. The underlying reasons are more nuanced. One important factor is the asymmetry observed in the SET/RESET I-V characteristics of measured samples. This asymmetry is not fully captured by adsorption energies alone, suggesting that additional effects are at play during both switching processes. For instance, the Au electrode in MoS2/Au heterostructures induces strain on the MoS2 layer, leading to charge transfer and modifications in the electronic structure [34], which can affect the energetics of both the SET and RESET processes.
S2 Details on adsorption energetics
Figure S3: Adsorption energies calculated with the inclusion of van der Waals forces using a nonlocal vdW-DF functional as implemented in VASP.
Figure S3 shows the variation in adsorption energies with the consideration of van der Waals (vdW) forces and without. In the majority of cases, vdW corrections yield adsorption energies that are nearly identical to those calculated without considering vdW interactions. For group 10 and 11 metals (Ni, Pd, Pt, Cu, Ag, Au), there is a consistent shift of approximately 2.36 eV when vdW corrections are included. Nevertheless, the trends in adsorption energy are retained, which is the focus of our current study.
Figure S4: Adsorption energies of the different TM adsorbates on MoSe2, WS2, WSe2 (blue) and the change in Bader charge on the adatom (red) from the neutral state where negative values correspond with charge gain and positive values correspond with charge loss.
Figure S5: Adsorption energies of the different TM adsorbates on MoSe2, WS2, WSe2 (blue) using the chemical potential derived from a bulk adsorbate and the Bader charge (red)
Figure S6: Comparison of the d-band centers of the adsorbates, referenced to the Fermi energy, with their respective adsorption energies on MoSe2. A d-band center of zero indicates alignment with the Fermi energy. The plots are categorized by specific d-block elements ((a) 3d, (b) 4d, (c) 4d, (d) all together) and color-coded by periodic groups. Linear regression lines and R2-values are provided as references to guide the eye.
Figure S7: Comparison of the d-band centers of the adsorbates, referenced to the Fermi energy, with their respective adsorption energies on WS2. A d-band center of zero indicates alignment with the Fermi energy. The plots are categorized by specific d-block elements ((a) 3d, (b) 4d, (c) 4d, (d) all together) and color-coded by periodic groups. Linear regression lines and R2-values are provided as references to guide the eye.
Figure S8: Comparison of the d-band centers of the adsorbates, referenced to the Fermi energy, with their respective adsorption energies on WSe2. A d-band center of zero indicates alignment with the Fermi energy. The plots are categorized by specific d-block elements ((a) 3d, (b) 4d, (c) 4d, (d) all together) and color-coded by periodic groups. Linear regression lines and R2-values are provided as references to guide the eye.
Figure S9: Heat map of d-orbital contributions of the different adatoms onto a sulfur-vacant MoS2 and selenium-vacant MoSe2
Figure S10: Heat map of d-orbital contributions of the different adatoms onto a sulfur-vacant WS2 and selenium-vacant WSe2
Table S1: Adsorbates Exhibiting or Not Exhibiting Metallic-like Behavior on TMD Surfaces
Observed Behavior Adsorbates
Exhibits Enhanced Metallic-like Behavior Ag, Au, Cd, Co, Cr, Cu, Fe
Hf, Hg, Mn, Mo, Ni, Os, Pd
Pt Ru, Sc, Ta, Ti, W, Y, Zn
Does Not Exhibit Enhanced Metallic-like Behavior Rh, V, Zr, Nb, Re, Tc, Ir
S3 Additional SISSO details
In the construction of $\Phi_{1}$, $\Phi_{2}$, and $\Phi_{3}$ we utilized basic algebraic operators given in Equation S1
$\hat{\mathbf{H}}^{(m)}\equiv\left\{I,+,-,\times,/,^{2},\sqrt{\text{ }},|-|\right\}$ (S1)
In addition, to prevent potential overfitting of data, we limit the maximal complexity of operators in a feature to 3.
Feature Name Abbreviation
1st Ionization Energy of adatom $IE_{\text{ads}}$
Electronegativity of adatom $X_{\text{ads}}$
Covalent Radius of Adatom $R_{c}^{\text{ads}}$
Weighted Workfunction of Adatom $\text{WF}_{\text{ads}}$
Electronegativity of cation $X_{\text{cat}}$
Covalent Radius of cation $R_{c}^{\text{cat}}$
Weighted Workfunction of cation $\text{WF}_{\text{cat}}$
Electronegativity of anion $X_{\text{an}}$
Covalent Radius of anion $R_{c}^{\text{an}}$
Weighted Workfunction of anion $\text{WF}_{\text{an}}$
Table S2: Primary features used in the SISSO-generated descriptor for adsorption energy. The table includes the electronegativity, covalent radius, and 1st ionization energy of adatoms, cations, and anions, as well as their weighted work functions. The electronegativity, covalent radius, and ionization energy values were obtained from [72], while the weighted work function values were obtained from [76]